Main INDEX
Monthly INDEX
PREV
NEXT
User name J. Singh
Log entry time 05:22:51 on May22,2003
Entry number 103261
keyword=septum dump, summary, part 1
since around 3ish, MCC has been doing HARP scans. the beam positions were all over the place, see following entries. about 0440, the septum dumped and started to warm up. since it looks like it will recover fairly quickly, we would call this a dump w/o quench. the temperatures seem to be easing back down, but the LHe level is decreasing which seems to make sense b/c of boil off.
we noticed that the read current was 0A about 5 min after the dump occurred. the power supply was off and i set the set current back to 0. now we are waiting for the septum to cool back down and the liquid helium level to rise back up.
after the dump, i called MCC to inquire about what might have happened. they said that right before the dump occurred, they had downloaded new parameters to their quads. this affected the beam steering. they had to steer the beam back b/c it was way off.
we may have seen other signs of the beam steering in our T1/T3 rates as well as the beam position earlier. see following entries.
the plot is the zoomed in septum strip chart when the dump occurred. dark green is the septum current. blue & red are the coil and yoke temps. light green is the beam current. the others are the Liquid He levels and flows.
FIGURE 1
FIGURE 2